### v1

# PCDreamer: Point Cloud Completion Through Multi-view Diffusion Priors

Abstract
This paper presents PCDreamer, a novel method for point cloud completion. Traditional methods typically extract features from partial point clouds to predict missing regions, but the large solution space often leads to unsatisfactory results. More recent approaches have started to use images as extra guidance, effectively improving performance, but obtaining paired data of images and partial point clouds is challenging in practice. To overcome these limitations, we harness the relatively view-consistent multi-view diffusion priors within large models, to generate novel views of the desired shape. The resulting image set encodes both global and local shape cues, which is especially beneficial for shape completion. To fully exploit the priors, we have designed a shape fusion module for producing an initial complete shape from multi-modality input (i.e., images and point clouds), and a follow-up shape consolidation module to obtain the final complete shape by discarding unreliable points introduced by the inconsistency from diffusion priors. Extensive experimental results demonstrate our superior performance, especially in recovering fine details.
1 Introduction
Point cloud completion plays a pivotal role in the advancement of 3D vision, serving as a crucial step in numerous applications, including autonomous driving [29, 17], robotics [4, 36], and augmented reality [10]. Incomplete point clouds, often resulting from scanning occlusions or limited sensor range, pose significant challenges. Although many encouraging approaches [49, 15, 44, 41, 48, 43, 45, 33, 55, 6, 56, 30, 40, 51] have been proposed, there is still considerable room for improvement regarding global completeness and local geometric details of resulting point clouds.
This paper addresses this task, focusing on single-view partial point clouds with self-occlusion [49, 15, 44, 45, 48, 43, 33, 40, 55, 6, 30]. As shown in Fig. 1(a) and (b), due to the single-view nature of the scanning process, the self-occlusion often leads to the loss of more than half of the 3D shape. Typical cases, like the lamp without the whole backside and the chair without most of the right side of the seat and legs, are shown in Fig. 1. The big missing region implies a vast solution space for completion.
Most existing methods [56, 30, 55, 6] rely solely on geometric information from partial point clouds to generate a complete shape in a coarse-to-fine manner. Although lacking the symmetric supporter of the chair back (Fig. 1(c)), the overall chair shape is completed with a certain amount of detail. However, the partial lamp without the critical top holder fails those methods, resulting in a random local guess. To address this, some approaches [51, 1, 41, 56] incorporate additional image information to improve completeness (Fig. 1(d), the chair), but obtaining paired image data is challenging in practice.
Motivated by the success of diffusion-based image generation [25, 23], we observe that the relatively view-consistent multi-view images generated by large diffusion models provide rich global and local shape cues, which are particularly advantageous for shape completion. For instance, when conditioned on partial point clouds, the generated images capture symmetric elements, such as the back of the chair (Fig. 1(e) top) and the top holder of the lamp (Fig. 1(e) bottom), enabling the full shape completion with all relevant features.
To harness the diffusion priors, we propose PCDreamer, a novel algorithm composed of three key components for high-quality point cloud completion. We, for the first time, introduce diffusion priors in this task. Specifically, in the first module, given the single-view partial point cloud, we utilize a large multi-view diffusion model to ‘dream’ a set of realistic images for free. These multi-view images depict the desired shape, serving as the fuel for the completion. The second module employs the attention mechanism to effectively fuse both the images and partial point cloud observation, leading to an initial full shape that preserves both the global structure and local features. Additionally, to address the inconsistency introduced by the multi-view images, the final module incorporates a confidence-guided shape consolidator, which produces the final complete shape with high fidelity. Extensive evaluation and experiments demonstrate the superior performance of our approach.
2 Related Work
Partial point cloud-based shape completion. Point cloud completion has traditionally relied on partial point clouds, often using intermediate 3D representations processed by 3D convolution, which suffer from resolution limits and high computational costs [8, 11, 36, 32]. Point-based networks like PointNet [26, 27] emerged to process 3D coordinates directly, with PCN [49] being the first to use an end-to-end point-based approach, generating points through coarse-to-fine folding. Subsequent encoder-decoder methods [37, 22, 43, 40] improved performance.
Recent attention-based approaches, particularly those using vector attention mechanisms [53], have further advanced the field. SnowflakeNet [43] and CRA-PCN [30] introduced novel transformer-based components for improved detail generation. SeedFormer [55] and AnchorFormer [6] both predict shape representations that preserve local information and integrate these representations into the generation process to restore high-fidelity details. Methods like PoinTr [48] use the partial point cloud as a condition, leveraging Transformer architectures to predict missing points. Despite the improvements, Transformer-based approaches still struggle with severely incomplete point clouds due to limited supplementary information.
Partial point cloud & Images-based shape completion. Given the limited geometric information in partial point clouds, other approaches incorporate additional data like RGB images and depth images to enhance completion performance. MVCN [14] uses a conditional GAN for completion in the 2D domain by leveraging rendered depth images from 3D shapes, but this results in spatial information loss. ViPC [51] addresses this by combining partial point clouds with single-view RGB images, though RGB images alone provide insufficient spatial detail. Similarly, Cross-PCC [41] employs single-view RGB images and 2D depth images from complete objects, achieving unsupervised point cloud completion without requiring full 3D point clouds. SVDFormer [56] utilizes multi-view depth images from partial point clouds to generate global shapes and produce new points by incorporating learned shape priors and geometric information. However, these depth images, though containing some spatial information, lack comprehensive geometric priors.
Multi-view images generation. Early works [35, 31, 24] employed specialized attention mechanisms to facilitate cross-view interactions, but often struggled with achieving fine detail consistency and incurred high computational costs. Wonder3D [25] tackled these limitations using a diffusion model equipped with cross-domain attention to generate both multi-view images and normal images. More recent approaches [19, 16, 13, 39, 28] have enhanced generation quality by incorporating novel attention mechanisms and generating multiple modalities.
Video generation presents even greater challenges due to the need for maintaining temporal consistency. Unlike previous methods [54, 12, 3, 9, 18, 52], Stable Video Diffusion [2] introduces a strategy of fine-tuning the entire video diffusion model, rather than focusing solely on temporal layers. This approach provides strong multi-view 3D priors, significantly improving cross-view consistency and generalization, and achieving superior results compared to image-based methods.
3 Method
Given a partial, non-uniform point cloud , our goal is to produce a dense, uniform, and complete point cloud that preserves ac consistent global shape while retaining detailed local features. Fig. 2 provides an overview of our method consisting of three core components, i.e., multi-view image generation, multi-modality shape fusion, and shape consolidation. In the following, we elaborate on the details.
3.1 Multi-view Image Generation
Given a partial point cloud observation, we aim to ‘dream’ out the full shape. To this end, we resort to the large multi-view diffusion models. There are a few choices for multi-view generation. For instance, RichDreamer [28] can conditionally generate multi-view depth images. However, the resulting images are of low resolution and poor consistency. To the best of our knowledge, no existing large model is capable of fulfilling this specific purpose. We thus designed a pipeline composed of a few large models in place (Fig. 2, bottom left).
Since most of the multi-view diffusion models take as input an RGB image, we first employ ControlNet [50] to translate the single-view depth image into an RGB image. We then pass the RGB image to a large diffusion model, for example, Wonder3D [25], which produces six posed multi-view images of the input shape. Note that, our method is not restricted to specific diffusion model, and we have experimented with Wonder3D and Stable Video Diffusion [2] (see Tabs. 1, 2, 5 and 4), as a proof of concept. Finally, DepthAnything[46] is exploited to generate corresponding depth images for each view, denoted as and for the depth image and the camera pose, respectively. The purpose of having depth images is to extract more accurate shape information from the 2.5D images. See our ablation study (Sec. 4.3) on the visual signal selection.
3.2 Multi-modality Shape Fusion
Having the multi-view depth images of the imagined shape, next, the key is to effectively fuse them into a full 3D shape that respects the golden partial observation, while accepting the missing regions from multi-view images. We thus designed a multi-modality shape fusion module with two specialized encoders and a fuser.
Partial Point Cloud Encoder. Due to the nature of single-view scanning, the partial observation is a golden standard for the desired shape, containing global and local geometric priors. We thus adopted a patch-based Transformer encoder to effectively extract geometric features, as shown in Fig. 3 (a). Formally, we first select seed points from the partial point cloud to divide it into patches , based on a fixed radius . Each patch goes into a DGCNN [38] encoder to extract patch-level features, while the center point (i.e., the seed point) of each patch is transformed into a positional encoding via the sinusoidal function. Finally, the Transformer encoder takes as input the concatenated features and outputs a unified feature of the partial point cloud after average pooling, defined as:
Multi-view Image Encoder. Given the posed multi-view depth images, a natural way is to fuse them into 3D and extract features from the fused 3D shape directly. We have proved in the ablation study in Sec. 4.3 that the resulting shape is irregular and noisy, because of the inherent inconsistency of the multi-view generation. To accurately extract global and local shape cues from the images, particularly for the missing regions, while mitigating the impact of inconsistencies, we have designed a patch-based image encoder, as shown in Fig. 3 (b). Specifically, each image goes through a pre-trained ResNet backbone to obtain the image feature, while its corresponding camera pose is further encoded via a two-layer MLP to create the positional encoding. Similarly, the Transformer takes as input the concatenated feature and produces the final multi-view image fusion feature after average pooling, defined as:
Mutli-modality Fuser. In order to effectively exploit the complementary information in both modalities, we resort to the attention mechanism to fuse the aforementioned features. Formally, we treat the point cloud feature as the query, while the the image feature as the key and value, defined as:
where the weights are learnable parameters. Then the attention is calculated as follows:
A final three-layer MLP converts the fused feature into an initial coarse point cloud .
3.3 Shape Consolidation
Due to the inherent inconsistencies in multi-view depth images and limitations in network fitting capabilities, the predicted coarse point cloud tends to be noisy and with localized holes. To address these issues, we further propose a confidence-based shape consolidator to produce the final dense completed point cloud.
Point Filtering. As shown in Fig. 1 bottom right, we assign confidence scores to each point within by examining 1) the consistency between each point and the multi-view feature and 2) the intra-agreement between points. Specifically, a point coordinate first goes through an MLP layer to obtain a corresponding feature, which is then concatenated with , as . We then calculate the average self-attention scores among all the concatenated features:
We empirically divided into high and low confidence sets ( vs. ) based on the calculated confidence score, and take the first set as a filtered intermediate point cloud that is significantly regular and compact.
Methods | Plane | Cabinet | Car | Chair | Lamp | Couch | Table | Boat | CD-Avg | DCD-Avg | F1 |
PCN [49] | 5.82 | 10.91 | 9.00 | 11.09 | 11.91 | 12.06 | 9.56 | 9.12 | 9.93 | - | 0.657 |
GRNet [44] | 8.51 | 12.84 | 10.17 | 12.13 | 11.83 | 12.64 | 11.16 | 9.84 | 11.14 | 0.765 | 0.563 |
PoinTr [48] | 4.31 | 9.23 | 7.60 | 8.35 | 8.27 | 9.38 | 7.97 | 7.01 | 7.76 | 0.562 | 0.810 |
PMP-Net++ [40] | 4.80 | 11.70 | 8.96 | 8.93 | 7.04 | 12.25 | 8.57 | 6.88 | 8.64 | 0.714 | 0.683 |
SnowFlakeNet [43] | 3.95 | 8.82 | 7.52 | 7.48 | 6.34 | 8.88 | 6.61 | 6.10 | 6.96 | 0.567 | 0.828 |
SeedFormer [55] | 3.87 | 9.05 | 7.53 | 7.63 | 6.12 | 9.26 | 6.61 | 6.03 | 7.01 | 0.573 | 0.824 |
AnchorFormer [6] | 3.62 | 8.79 | 7.20 | 7.12 | 6.18 | 8.71 | 6.50 | 6.01 | 6.77 | 0.539 | 0.841 |
HyperCD [21] | 3.87 | 9.05 | 7.40 | 7.50 | 5.95 | 9.25 | 6.50 | 5.89 | 6.93 | - | - |
SVDFormer [56] | 3.68 | 8.73 | 7.10 | 6.95 | 5.64 | 8.58 | 6.26 | 5.91 | 6.61 | 0.534 | 0.848 |
CRA-PCN [30] | 3.62 | 8.77 | 7.00 | 6.92 | 5.46 | 8.59 | 6.27 | 5.86 | 6.56 | - | - |
Ours(SVD-MV) | 3.51 | 8.62 | 6.92 | 6.91 | 5.66 | 8.31 | 6.27 | 5.90 | 6.52 | 0.531 | 0.856 |
Ours(Wonder3D) | 3.52 | 8.72 | 6.89 | 6.71 | 5.64 | 8.32 | 6.24 | 5.84 | 6.49 | 0.518 | 0.859 |
Final Consolidation. Firstly, we simply merge and to become a denser point cloud, then apply farthest point sampling (FPS) to obtain a new point set . We finally upsample to produce the dense and uniformly distributed complete point cloud.
Specifically, we use an MLP to extract per-point shape features () from and point-wise concatenate them with the , resulting in a combine feature of dimension . The fused features are then projected to a high-dimensional space to generate a set of offsets (i.e., times the number of points), defined as:
where reshapes to . We then map it to by a one-layer MLP. Finally, for each point in , we add the offsets from to produce the final completion result .
Network Training
Our whole network is trained end-to-end with the Chamfer distance loss. Specifically, we measure the discrepancy between the ground truth point cloud and the coarse () and final () point clouds, respectively. Inspired by [47], we exploit the extra hyperbolic Chamfer distance (HyperCD) to combat noise and outliers, defined as:
and
where is the Euclidean distance between point and . The final loss is:
Methods | Table | Chair | Plane | Car | Sofa | Dishwasher | Mailbox | Microwave | Pillow | Remote | CD-Avg | DCD-Avg | F1 |
PCN [49] | 1.80 | 1.59 | 0.81 | 1.44 | 1.52 | 1.50 | 1.73 | 1.83 | 1.33 | 0.78 | 2.064 | 0.667 | 0.194 |
GRNet [44] | 1.72 | 1.54 | 0.82 | 1.56 | 1.61 | 1.77 | 1.39 | 1.91 | 1.39 | 0.79 | 1.801 | 0.696 | 0.243 |
PoinTr [48] | 1.62 | 1.23 | 0.86 | 1.13 | 1.00 | 1.00 | 0.84 | 1.22 | 0.73 | 0.51 | 1.062 | 0.643 | 0.387 |
AnchorFormer [6] | 1.12 | 0.81 | 0.43 | 1.11 | 0.93 | 1.21 | 0.81 | 1.22 | 0.71 | 0.36 | 1.182 | 0.670 | 0.328 |
SVDFormer [56] | 1.03 | 0.80 | 0.44 | 1.00 | 0.88 | 1.00 | 0.76 | 1.27 | 0.71 | 0.34 | 0.938 | 0.621 | 0.447 |
CRA-PCN [30] | 1.58 | 1.08 | 0.58 | 1.24 | 1.19 | 1.31 | 1.18 | 1.31 | 0.86 | 0.45 | 1.300 | - | - |
Ours(SVD-MV) | 0.99 | 0.75 | 0.41 | 0.99 | 0.88 | 0.94 | 0.61 | 1.18 | 0.65 | 0.34 | 0.907 | 0.559 | 0.408 |
Ours(Wonder3D) | 0.98 | 0.71 | 0.40 | 0.96 | 0.85 | 0.98 | 0.60 | 1.21 | 0.62 | 0.32 | 0.895 | 0.569 | 0.418 |
4 Experiments
4.1 Dataset
PCN dataset. The PCN dataset [49] is a subset of the ShapeNet dataset [5], containing shapes from eight categories. In this study, we constructed our dataset based on PCN. Specifically, given the mesh data of a complete object, we selected a fixed viewpoint (the same as the Wonder3D viewpoint) along the positive z-axis on its unit sphere to obtain a single-view depth image and generated a partial point cloud with 2,048 points through a back-projection method. Additionally, multi-view images were produced from the single-view depth image using the multi-view image generation module (Sec. 3.1). We uniformly sampled 16,384 points from the mesh surface to obtain the corresponding ground truth point cloud for each object. In total, we collected 30974 samples and split the dataset into training, validation, and test sets, the ratio is the same as [49].
ShapeNet-55 dataset. Unlike the PCN dataset [49], ShapeNet-55 [48] includes shapes from 55 categories. We apply the same construction method as aforementioned to create a dataset.
Evaluation metrics. Following recent works [55, 56], we adopt Chamfer Distance (CD), Density-aware CD (DCD) [42], and F1-Score as evaluation metrics, and report the version of CD for PCN dataset and the version of CD for Shapenet-55 dataset.
Implementation. Our model is trained for 300 epochs on both datasets using four NVIDIA 4090 GPUs with a batch size of 24. We employ the Adam optimizer with an initial learning rate of 1e-4. It is decayed by a factor of 0.7 every 40 epochs on the PCN dataset [49] and by a factor of 0.98 every 2 epochs on the ShapeNet-55 dataset [48]. In all of our experiments, we set K to 128 and r to 0.2. For ControlNet, we set the control strength to 1.0 and the guidance scale to 5.0. For Stable Video Diffusion (SVD-MV), we set the elevation angle to 5 degrees, leveraging its strong generalization capability, while keeping all other parameters at their default values. The resulting video contained 21 frames, from which we selected 6 fixed frames as the multi-view images. Finally, we normalized the generated depth image values to the range of 0 to 1.
4.2 Comparisons
To demonstrate the effectiveness of our method, we compare it with SoTA methods, including PCN [49], GRNet [44], PoinTr [48], PMP-Net++ [40], SnowFlakeNet [43], SeedFormer [55], AnchorFormer [6], HyperCD [21], SVDFormer [56], and CRA-PCN [30]. We retrain their networks from scratch on our datasets with their default configurations. The quantitative and qualitative results are presented in the following.
Results on PCN Dataset. The quantitative results in Tab. 1 demonstrate that our method outperforms all previous approaches across nearly all categories while achieving the best average performance. An exception is the Lamp, where CRA-PCN achieves slightly better performance. The inconsistency in the multi-view image generation module, especially when the lamp’s support structure has a folding line or curved shape, leads to lower accuracy in our method. A visual comparison is presented in Fig. 4, where the visual results in several selected categories (e.g., Plane, Chair, and Boat) show that our method can predict the missing parts more accurately due to the additional shape cues provided by the generated multi-view images. Particularly, for the Airplane category, all other methods fail to complete the missing symmetric parts (e.g., the engines), while our results are more consistent with the ground truth.
Results on the ShapeNet-55 Dataset. Following existing work [56], we also compare our method against some competitors [49, 44, 48, 6, 56, 30] on the ShapeNet-55 dataset. We did not compare with SeedFormer-based methods [55, 21] because we have observed occasional numerical failure during their training caused by the occurrence of NaN values. As shown in Tab. 2, we selected 10 representative classes from a total of 55 categories to showcase the Chamfer Distance (CD). Our method consistently outperforms other state-of-the-art approaches across all categories, with a particularly notable improvement of 0.16 in the CD metric for the Mailbox category. Although the first five classes contain significantly more samples than the latter five, we have obtained comparable results across the ten categories, which demonstrates the generalization ability and robustness of our method on diverse types of data. The visual comparison in Fig. 5 also clearly illustrates the superior performance on more complex shapes with finer details.
Comparison with 3D model generation.
We further compare our approach with recent methods [23, 20] that are pre-trained on large-scale datasets and leverage powerful diffusion models to generate complete 3D models from a single-view RGB image, thereby enabling effective shape completion as well. For a fair comparison, we employ ControlNet to generate an RGB image from a single-view depth image, which is then used as input to their pretrained models to predict the complete 3D shape. Finally, the point cloud is extracted using Poisson disk sampling. As shown in Fig. 6, although One-2-3-45++ [23] and Crafsman [20] can generate overall regular shapes, they fail to capture local geometric features, e.g., the back of the chair and the symmetric engines of the airplane, due to the limited cues from the single-view image. Instead, our methods can faithfully recover a full shape with fine details based on the cues from the multi-view images. More comparison results can be found in the supplementary.
4.3 Ablation study
We have extensively ablated our core design choices, and report both the qualitative and quantitative results on the PCN dataset in the following.
Depth images | Partial input | CD-Avg | DCD-Avg | F1 |
✓ | 17.69 | 0.823 | 0.401 | |
✓ | 6.91 | 0.534 | 0.840 | |
✓ | ✓ | 6.49 | 0.518 | 0.858 |
RGB images | depth images | CD-Avg | DCD-Avg | F1 |
✓ | 6.63 | 0.530 | 0.851 | |
✓ | 6.49 | 0.518 | 0.858 | |
✓ | ✓ | 6.59 | 0.535 | 0.851 |
Is depth image a good choice? From the multi-view image generation module, we obtained both the depth and RGB images. The former contains 2.5D information of the underlying geometry, while the latter presents more vivid appearance cues. In this test, we use either the RGB image, the depth image, or their combination as our multi-view signal in the later modules in our pipeline, and present the statistical results in Tab. 4. As can be seen, the depth-only option outperforms others across multiple metrics. The main reason is that the RGB images generated by ControlNet occasionally contain unrealistic artifacts or inconsistencies, which can adversely affect the completion results.
Does multi-modality fusion play an important role? To evaluate the effectiveness of our multi-modality fusion strategy, we split the network architecture into two branches. The first branch generates a complete point cloud directly from multi-view images, i.e., after getting from , we use an MLP to map it to a coarse point cloud, which is then refined by the decoder module to produce the final result. The second branch operates solely on the partial point cloud by removing the multi-view image module from our pipeline. Quantitative results of Tab. 3 demonstrate that the multi-modality feature fusion strategy achieves the best performance. As shown in Fig. 7, incorporating additional information from multi-view images allows the model to better capture both the global shape and local structures, such as the backrest of a chair and the supporting regions of a table.
Is confidence filtering helpful? To evaluate the impact of our proposed confidence-based filtering operation, we conducted an ablation study by removing it from our method. As shown in Fig. 8, the red box in (b) highlights noisy points that are assigned low confidence scores by our method. By applying the filtering strategy, as shown in (c), these noisy points are effectively removed, resulting in a consolidated and improved completion result. Quantitative results in Tab. 5 further demonstrate the effectiveness of the confidence filtering operation.
Methods | CD-Avg | DCD-Avg | F1 |
w/o confidence filtering | 6.76 | 0.525 | 0.845 |
Ours | 6.49 | 0.518 | 0.858 |
4.3.1 Discussion
Number of views. The guidance provided by multi-view depth images is crucial for point cloud completion, and the number of views is a key hyperparameter that significantly affects the results. To investigate, we have conducted a series of experiments. First, the SVD-MV model generates 21 frames for each object. We use this as a candidate set and randomly select views from it to train the model separately. As shown in Tab. 6, fewer views result in limited coverage of the object’s overall shape. As increases, the entire shape becomes fully covered, and further increment of only leads to larger network parameters without significantly improving performance.
Hyperparameters of Filtering. The filtering strategy can effectively reduce noise in the coarse point cloud, thereby improving the accuracy of completion. However, the number of points filtered needs to be carefully balanced. As shown in Tab. 5, filtering too many points can lead to a loss of shape features, affecting the final accuracy, while filtering too few points may retain noise, also impacting precision. To address this, we conducted experiments with different filtering levels and selected the optimal setting.
View number | CD-Avg | DCD-Avg | F1 |
2 | 6.65 | 0.542 | 0.842 |
4 | 6.64 | 0.539 | 0.848 |
6 | 6.52 | 0.531 | 0.856 |
8 | 6.62 | 0.536 | 0.846 |
Filtered points | CD-Avg | DCD-Avg | F1 |
12% | 6.92 | 0.533 | 0.840 |
25% | 6.49 | 0.518 | 0.858 |
50% | 6.47 | 0.530 | 0.857 |
Limitation. Multi-view depth images can guide point cloud completion, but there are limitations in generating these images. As shown in the Fig. 9, the slender structure of the earphone exhibits significant inconsistencies across different generated views, resulting in poor point cloud completion when used as guidance. In cases like this, where the information visible in a single view is extremely limited, the completion results are often unsatisfactory.
5 Conclusion
In this paper, we have presented a novel approach for point cloud completion. We, for the first time, introduce diffusion priors in this task to provide imagined full-shape cues for free. From this proof of concept, we believe that diffusion priors can be useful for other point cloud processing tasks, and we hope PCDreamer can inspire future research in this direction.
Supplementary Material
In this supplementary, we present more visual results on the PCN [49] and ShapeNet-55 [48] datasets. We also present additional visual examples, comparing our approach against more single-view 3D generation methods. Besides, to further demonstrate the generalization ability of our method in real-world scans, we also conduct experiments on the ScanNet dataset [7], which contains partial point clouds extracted from RGB-D scans.
Additional results on Real-world Scans.
We evaluated our method on real-world scanned data from the ScanNet dataset [7] to assess its performance. Specifically, we selected depth images of individual objects and their corresponding partial point clouds from ScanNet. Using our multi-view generation module, we obtained multi-view depth images. These were then input into our model, which was pre-trained on the PCN dataset, to produce the completion results. As shown in Fig. 1, our method yields reasonable and high-quality results on real scanned data, demonstrating its effectiveness and robustness in handling real-world inputs.
Visual comparison.
In Fig. 2, we present additional comparisons with other 3D model generation methods [39, 34]. Moreover, we provide further examples from the PCN dataset for comprehensive evaluation. As illustrated in the figure, the lack of extra views, the randomness in the generation process, and the inherent inconsistencies in multi-view images often prevent single-view 3D model generation methods from effectively capturing geometric details. Consequently, these models may produce shapes that, while plausible, exhibit significant deviations from the ground truth, such as variations in the width of the sofa. Therefore, multi-view images are more appropriate as auxiliary cues for point cloud completion, although directly utilizing them for point cloud completion typically yields suboptimal results.
Additional results on PCN and ShapeNet-55 dataset.
Figs. 3 and 4, each presents an extra six results from the PCN and the ShapeNet-55 datasets, respectively. For both datasets, the first three cases in the first three rows possess high-quality and consistent multi-view depth images, whereas depth images with lower quality and less consistency are generated for the last three examples. Specifically, the high randomness of the generation process, combined with limitations of the initial view, can result in multi-view depth images exhibiting inconsistencies across views (e.g., the lid of the Dustbin), missing geometric details(e.g., the Helmet), and the presence of noisy regions (e.g., the leg of the Table). Furthermore, shapes imagined by the generative model may exhibit scale discrepancies relative to their real-world counterparts (e.g., the Sofa and the Table). In addition, the depth estimation process may also occasionally produce suboptimal outcomes (e.g., the Dishwasher). However, the proposed confidence-based shape consolidator effectively addresses these issues by eliminating unreliable points caused by inconsistencies in diffusion priors. As a result, our models produce accurate and reasonable outputs, as demonstrated in the second and fourth columns of the last three examples.